A vase was originally sold for RM27000. The price was marked up by 25% and finally marked down by 25%. What is the final price?
Found 2 solutions by Alan3354, Theo:
Answer by Alan3354(69443)

A vase was originally sold for RM27000. The price was marked up by 25%
---
27000 + 27000*0.25 = 33750
-----------------
and finally marked down by 25%. What is the final price?
33750 - 33750*0.25 = 25312.50

Answer by Theo(13342)

the original price was 27000.

this price was marked up by 25% and then the new price was marked down by 25% to get the final price.

the mark up on the original price was 25% of the original price which makes it .25 * 27000 = 6750.

that mark up was added to the original price to make the marked price equal to 27000 + 6750 = 33750.

this marked up price of 33750 was marked down 25%.

25% of 33750 = 8437.5

that mark down was subtracted from 33750 to make the marked down price equal to 33750 - 8437.5 = 25312.5

that's the final price.

note that 27000 + .25 * 27000 is equal to (1 + .25) * 27000 which is equal to 1.25 * 27000.

note that 33750 - .25 * 33750 is equal to (1 - .25) * 33750 which is equal to .75 * 33750.

the marked up price is equal to 1.25 * 27000 which is equal to 33750

the marked down price is equal to .75 * 33750 which is equal to 2531.5.

you start with the original price of 27000 and multiply it by 1.25 to get the marked up price and then multiply it by .75 to get the marked down price.

27000 * 1.25 * .75 = 25312.5

that's the final price which is the original price marked up by 25% which is then marked down by 25%.

the mark up is from 27000.

the mark down is from 33750.